Given x + iy form, write in polar form: negative radical 5 + i
Answer by stanbon(48564)

Given x + iy form, write in polar form: negative radical 5 + i
-sqrt(5) + i
r=sqrt[(sqrt(5))^2 + 1^2] = sqrt[5+1] = sqrt(6)
theta = arctan(1/(-sqrt(5))=155.91 degrees
polar form = (sqrt6)[sin(155.91)+icos(155.91)]
Cheers,
Stan H.